OpenStudy (anonymous):

how do you solve system of equations? 3x+5y=9 2x-y=-7

OpenStudy (anonymous):

why did you multiply the first eqn by 2 and the second by 3?

OpenStudy (anonymous):

I mean negative 3?

OpenStudy (anonymous):

Or let someone else explain:)

OpenStudy (anonymous):

don't we have to put y by itself and then solve? why multiply by 5? I don't quite get the concept

OpenStudy (aripotta):

there's a number of ways we can solve this. substitution, elimination, graphing, or matrix.

OpenStudy (aripotta):

which do you prefer?

OpenStudy (anonymous):

from the second equation we have 2x=y-7 or x=(y-7)/2 ok no we go to the first equation (3/2)*(y-7)+5y=9, =>(3/2)y-21/2+5y=9 (3/2)y+5y=9+(21/2) =>6.5y=9.5 or y=3 and from x=(y-7)/2 we have x=-1/2

OpenStudy (anonymous):

hmm substituion I guess.. Ximcan you please explain how you did that?

OpenStudy (anonymous):

Ximki*

OpenStudy (aripotta):

alright. solve for y in the second equation in terms of x. what do you get?

OpenStudy (anonymous):

I have to set x to 0 right?

OpenStudy (anonymous):

ok Aisha wait a minute I will be more clear

OpenStudy (anonymous):

thanks :)

OpenStudy (aripotta):

no, you have to rearrange the equation so that it's just y on the left side

OpenStudy (anonymous):

oh ok so I got -y=-2x-7

OpenStudy (aripotta):

right, buy you gotta divide both sides by -1

OpenStudy (anonymous):

y=2x+7

OpenStudy (aripotta):

but* to make y positive

OpenStudy (aripotta):

yea. so substitute that into 3x+5y=9

OpenStudy (aripotta):

then solve for x

OpenStudy (anonymous):

it will be 3x+5(2x+7)=9?

OpenStudy (aripotta):

now distribute

OpenStudy (anonymous):

x=-26/13

OpenStudy (aripotta):

yea, which is simplified to...?

OpenStudy (anonymous):

-2

OpenStudy (aripotta):

yes. now substitute that into either of the equations to solve for y

OpenStudy (anonymous):

is it (-2,3)?

OpenStudy (aripotta):

yes :)

OpenStudy (anonymous):

yay thanks a lot

OpenStudy (anonymous):

you really helped me :) the finals answer is -2,3 and I don't have to graph it roght?

OpenStudy (aripotta):

nah

OpenStudy (anonymous):

Thanks :)

OpenStudy (aripotta):

graphing would be a way to solve for it. but you don't have to plot the answer if you solved it a different way

OpenStudy (aripotta):

no problem :)

OpenStudy (anonymous):

oh okay gotcha :)

OpenStudy (anonymous):

sorry for late :P